The formula S= C(1+r)^t models inflation, where C the value today, r the annual inflation rate, and S the inflated value years from now. Use this formula to solve the following exercise.If the inflation rate is 5% , how much will a house now worth $760,000 be worth in 4 years?
Answer by checkley79(3341)

S= C(1+r)^t
S=760000(1+.05)^4
S=760000(1.05)^4
S=760000*1.2155
S=$923,780 ans.